At what rate of compound interest will a sum of money become 25/16 times of itself in 2 years?

Grade:10

2 Answers

Arun
3 years ago
Dear student
 
let the money = x
 
hence
 
25x/16 = x ( 1 + r/100)2
 
5/4 = 1 + r/ 100
¼ = r/100
 
r = 25%
 
Regards
Vikas TU
3 years ago
Let the sum be P 
ATQ = (25/16)P
25P/16 = P(1+R/100)^2
(5/4)^2 = (100+R)^2/(100)^2
500 = 400 + 4R 
R = 25% Per annum